OpenStudy (anonymous):

simplify rational expression... 2n/n-4 n= 6/x+3 x= and 3b/b(b+9) b=

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ 2n}{n} -4 }\) like this?

OpenStudy (anonymous):

the n-4 is together under the 2n

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{ 2n}{n-4} }\)

OpenStudy (solomonzelman):

Is that equal to something, or just that fraction to simplify?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it says to simplify the rational expression

OpenStudy (solomonzelman):

Well, I don't really see a way to "simplify" this ... you can do this ... \(\color{#000000 }{ \displaystyle \frac{ 2n}{n-4}= \frac{ 2n-8+8}{n-4} = \frac{ 2n-8}{n-4}+ \frac{ 8}{n-4}= \frac{ 2(n-4)}{n-4}+ \frac{ 8}{n-4}}\) and then cancel the (n-4) on top and bottom

OpenStudy (anonymous):

okay what about the other two

OpenStudy (anonymous):

6/x+3 and 3b/b(b+9) anyone please?

OpenStudy (fibonaccichick666):

the reason you (probably) are not getting answers is due to the questions a and b making minimal sense.

OpenStudy (anonymous):

i think that's the issue.. I'm trying to help my cousin with her homework but neither of us are understanding.

OpenStudy (fibonaccichick666):

You stated that there is no equals sign and that you are supposed to be simplifying the fraction, but in a and b, most would consider those simplified

OpenStudy (fibonaccichick666):

can you take a picture?

OpenStudy (anonymous):

i can

OpenStudy (fibonaccichick666):

I can then look at the context and see if it is poorly worded or if there is something missing

OpenStudy (anonymous):

I'm trying to post it give me a second

OpenStudy (fibonaccichick666):

okie

OpenStudy (anonymous):

How do I post from a cell phone?

OpenStudy (anonymous):

Or do I have to do it from the laptop? I have open study open on both

OpenStudy (anonymous):

It says "simplifying rational expressions find the excluded value(s) for each rational expression"

OpenStudy (anonymous):

Then states the questions 2n/n-4 N= And then the other two

OpenStudy (fibonaccichick666):

I think you can upload a pic with a cell

OpenStudy (fibonaccichick666):

OH

OpenStudy (fibonaccichick666):

Wow, ok this is poorly worded. The problem wants you to find the values that make the fraction undefined/indeterminate. IE denominator is equal to zero

OpenStudy (fibonaccichick666):

that's what an excluded value is

OpenStudy (anonymous):

Okay

OpenStudy (fibonaccichick666):

so in the first one, what will x have to equal for the bottom to equal zero?

OpenStudy (fibonaccichick666):

sorry, n*

OpenStudy (fibonaccichick666):

can you get it from here?